# Extending the blended generalized extreme value distribution

### v2

@@ 1. Introduction
The generalized extreme value (GEV) distribution can be derived as the limit distribution of block maxima (from any distribution for which such a limit exists) for sufficiently large blocks and numerous samples (the Fisher–Tippett–Gnedenko theorem) [1]. However, in practice the GEV distribution is fitted to finite data sets, such as observations of temperature, precipitation, streamflow, and sea level (among many other applications to problems in fields ranging from finance to biology), to estimate the likelihood of future extreme events [2, 3]. In this context, some properties of the GEV distribution complicate inference.
@@
$F_{\text{GEV}}(s)$ is the GEV distribution at $s=\frac{x-\mu}{\sigma}$ with the chosen parameters $\xi<0,\mu,\sigma$. $F_{\text{Gumbel}}(\tilde{s})$ is the Gumbel distribution with location and scale parameters $\tilde{\mu},\tilde{\sigma}$ used to standardize $\tilde{s}=\frac{x-\tilde{\mu}}{\tilde{\sigma}}$. $\tilde{\mu},\tilde{\sigma}$ are chosen such that $F_{\text{Gumbel}}(\tilde{s}_{a})=F_{\text{GEV}}(s_{a})=a$ and $F_{\text{Gumbel}}(\tilde{s}_{b})=F_{\text{GEV}}(s_{b})=b$ for two user-specified quantiles $01>a>b>$ 0. In this case $a,b$ will typically be closer to 1 so modification of the Weibull GEV distribution is concentrated near its problematic upper bound, as opposed to the $\xi>0$ case where $a,b$ will typically be chosen to be closer to 0. $p(s)$ is a function that equals 0 for $s\geq s_{a}$, 1 for $s\leq s_{b}$, and, for intermediate values $s_{b}<s<s_{a},$ is equal to the cumulative distribution function of the beta distribution $F_{\text{beta}}(\frac{s-s_{a}}{s_{b}-s_{a}};\alpha,\beta)$ with user-specified positive shape parameters $\alpha,\beta$. The probability density function (PDF) of the bGEV is also formally the same as that given by [7] in the mixing region, while being equal to the GEV distribution PDF in the lower part where $p(s)=1$ and to the Gumbel distribution PDF in the upper part where $p(s)=0$. Finally, the limit of the bGEV distribution as $\xi$ approaches 0 from either direction is simply the Gumbel distribution.
@@ 2. Example simulations
Simulation case A involved drawing samples from a GEV distribution with $\xi=-0.2,\mu=0,\sigma=1$. This distribution has zero probability for $x\geq\mu-\sigma/\xi=5$. Over 100 realizations, the maximum-likelihood GEV fit to 100 samples drawn from this distribution has, in the median, an upper threshold (with zero probability mass above it) of 4.70, meaning that while the generating distribution does have an upper threshold (of 5), most of the GEV fits have an upper threshold that is too low, so that there is a possibility of a surprise value from this distribution that is above the maximum possible in the fitted distribution. In one realization, the fitted GEV had an upper threshold as low as 2.83, the 98.5 percentile of the generating distribution. By contrast, the bGEV (with $a=0.91,b=0.90,s=5$) was fitted with negative $\xi$, but by design had no hard upper threshold, obviating the problem of zero probability mass assigned to part of the generating distribution, although the bGEV could not capture the drop-off of the generating probability density to 0 at $x=5$.
Simulation case B involved drawing samples from a GEV distribution with $\xi=-0.2,\mu=0,\sigma=1$ but with an independent, normally distributed term with mean 0 and standard deviation 1 added to each sample, so that there is no hard threshold anymore; rather, the upper tail falls off like that of a normal distribution. The GEV fitted to 100 samples still indicated a negative shape parameter, giving a median upper threshold of 5.00, which was now qualitatively incorrect since the generating distribution actually had no upper limit. The bGEV’s shape, with no upper threshold, was qualitatively correct. (Figure 1) (The bGEV parameters $a,b,s$ could potentially be adjusted in order to improve quantitative fit to the upper tail of the sampled distribution, if this was important in a real application.)
Figure 1. GEV and bGEV fits to synthetic data (case B). The GEV and bGEV probability distribution functions shown are averaged over 100 realizations, fit to 100 independent samples each. All 10,000 samples are included in the histogram.
@@ 3. Example applications
Recent years have seen long-standing temperature records in many countries broken by unprecedented amounts. The state-of-the-art ERA5 reanalysis [15, 16] shows this pattern well, with most of the globe reaching new temperature records in the last 15 years. As a preliminary evaluation of the ability of GEV and bGEV distributions to fit this type of data, a sample was taken of 100 populated land grid cells spaced at least 10° apart, sufficient for good global coverage. For each chosen grid cell, annual maximum temperatures for 1940-2023, in units of K, were determined from the ERA5 hourly 2-m air temperature series, resulting in a time series of $n=84$ block maxima. Such temperature series are frequently fitted using the GEV distribution, as for example reflected in work by the World Weather Attribution collaboration in estimating the anthropogenic contribution to the probability of observed heat extremes occuring [17, 18, 19].
The GEV and bGEV distributions were compared in terms of one-year-ahead probabilistic forecasts, using negative log likelihood (NLL) of the actual value in the forecast as the skill metric [20, 5, 21] (this is closely related to cross-entropy and Kullback–Leibler divergence [22, 23]). The probabilistic forecasts were generated by fitting, for each length $s$ from 30 up to $n-1$, maximum-likelihood GEV and bGEV distributions for the first $s$ annual maxima, and using those to generate a forecast for year $s+1$ that could be compared to the observed value. (The bGEV was allowed to have a shape paramater of either sign, using the extension outlined above.) NLL was therefore summed over 5400 separate forecasts (54 years for each of 100 locations).
@@
Figure 4. Example of a new sea level record from Hilo, Hawai’i in 2020, showing the histogram of the 1970-2019 annual maximum sea level and the forecasts for 2020 based on the GEV and bGEV. Both the GEV and bGEV appropriately shift the forecast probability distribution upward relative to the historic one dues to global warming and have almost the same forecast mode, but the GEV distribution, with negative shape parameter, has zero probability mass past 0.9143 m, whereas the bGEV forecast has positive probability for the actual 0.916 m (vertical line).

### v3

@@ 1. Introduction
The generalized extreme value (GEV) distribution can be derived as the limit distribution of block maxima (from any distribution for which such a limit exists) for sufficiently large blocks and numerous samples (the Fisher–Tippett–Gnedenko theorem) [1]. However, in practice the GEV distribution is fitted to finite data sets, such as observations of temperature, precipitation, streamflow, and sea level (among many other applications to problems in fields ranging from finance to biology), to estimate the likelihood of future extreme events [2, 3]. “Engineering design is primarily concerned with the extremes”, which, in the context of buildings and civil infrastructure, requires accurate assessment of the frequency of the most impactful weather events and how this might change over the design lifetime [4]. In this context, some properties of the GEV distribution complicate inference.
@@
$F_{\text{GEV}}(s)$ is the GEV distribution at $s=\frac{x-\mu}{\sigma}$ with the chosen parameters $\xi<0,\mu,\sigma$. $F_{\text{Gumbel}}(\tilde{s})$ is the Gumbel distribution with location and scale parameters $\tilde{\mu},\tilde{\sigma}$ used to standardize $\tilde{s}=\frac{x-\tilde{\mu}}{\tilde{\sigma}}$. $\tilde{\mu},\tilde{\sigma}$ are chosen such that $F_{\text{Gumbel}}(\tilde{s}_{a})=F_{\text{GEV}}(s_{a})=a$ and $F_{\text{Gumbel}}(\tilde{s}_{b})=F_{\text{GEV}}(s_{b})=b$ for two user-specified quantiles $01>a>b>$ 0. In this case $a,b$ will typically be closer to 1 so modification of the Weibull GEV distribution is concentrated near its problematic upper bound, as opposed to the $\xi>0$ case where $a,b$ will typically be chosen to be closer to 0. $p(s)$ is a function that equals 0 for $s\geq s_{a}$, 1 for $s\leq s_{b}$, and, for intermediate values $s_{b}<s<s_{a},$ is equal to the cumulative distribution function of the beta distribution $F_{\text{beta}}(\frac{s-s_{a}}{s_{b}-s_{a}};\alpha,\beta)$ with user-specified positive shape parameters $\alpha,\beta$.
The PDF of the proposed bGEV is also formally the same as that given by [8] in the mixing region, while being equal to the GEV distribution PDF in the lower part where $p(s)=1$ and to the Gumbel distribution PDF in the upper part where $p(s)=0$. Finally, the limit of the bGEV distribution as $\xi$ approaches 0 from either direction is simply the Gumbel distribution. Unlike for the GEV, the mean and variance of the bGEV do not have simple analytic forms, although they can be computed numerically by integrating over the PDF. In practice, these bGEV properties are often close to those of the corresponding GEV, although higher (for negative $\xi$) since the bGEV is the same as the GEV at the lower quantiles but is unbounded above. For example, if $\xi=-0.3,\mu=0,\sigma=1$, the GEV mean and variance (to 8 decimal places) are 0.34176435 and 0.97846332 respectively, while the bGEV values with $\xi=-0.3,\mu=0,\sigma=1,a=0.95,b=0.8,\alpha=\beta=5$ are 0.35018832 and 1.02559938. The CDF of the bGEV also does not have a simple analytic inverse in the mixing region where $F_{\text{bGEV}}(s(x))$ is between $a$ and $b$, but quantiles in that range can be found numerically; for example, for $\xi=-0.3,\mu=0,\sigma=1,a=0.95,b=0.8,\alpha=\beta=5$, the 90th percentile of the bGEV is 1.61258469.
@@ 2. Example simulations
Simulation case A involved drawing samples from a GEV distribution with $\xi=-0.2,\mu=0,\sigma=1$. This distribution has zero probability for $x\geq\mu-\sigma/\xi=5$. Over 100 realizations, the maximum-likelihood GEV fit to 100 samples drawn from this distribution has, in the median, an upper threshold (with zero probability mass above it) of 4.70, meaning that while the generating distribution does have an upper threshold (of 5), most of the GEV fits have an upper threshold that is too low, so that there is a possibility of a surprise value from this distribution that is above the maximum possible in the fitted distribution. In one realization, the fitted GEV had an upper threshold as low as 2.83, the 98.5 percentile of the generating distribution. By contrast, the bGEV (with $a=0.91,b=0.90,\alpha=\beta=5$) was fitted with negative $\xi$, but by design had no hard upper threshold, obviating the problem of zero probability mass assigned to part of the generating distribution, although the bGEV could not capture the drop-off of the generating probability density to 0 at $x=5$.
Simulation case B involved drawing samples from a GEV distribution with $\xi=-0.2,\mu=0,\sigma=1$ but with an independent, normally distributed term with mean 0 and standard deviation 1 added to each sample, so that there is no hard threshold anymore; rather, the upper tail falls off like that of a normal distribution. The GEV fitted to 100 samples still indicated a negative shape parameter, giving a median upper threshold of 5.00, which was now qualitatively incorrect since the generating distribution actually had no upper limit. The bGEV’s shape, with no upper threshold, was qualitatively correct. (Figure 1) (The bGEV parameters $a,b,\alpha,\beta$ could potentially be adjusted in order to improve quantitative fit to the upper tail of the sampled distribution, if this was important in a real application.)
Figure 1. GEV and bGEV fits to synthetic data (case B). The GEV and bGEV probability density functions shown are averaged over 100 realizations, fit to 100 independent samples each. All 10,000 samples are included in the histogram.
@@ 3. Example applications
Recent years have seen long-standing temperature records in many countries broken by unprecedented amounts. The state-of-the-art ERA5 reanalysis [16, 17] shows this pattern well, with most of the globe reaching new temperature records in the last 15 years. As a preliminary evaluation of the ability of GEV and bGEV distributions to fit this type of data, a sample was taken of 100 populated land grid cells spaced at least 10° apart, sufficient for good global coverage. For each chosen grid cell, annual maximum temperatures for 1940-2023, in units of K, were determined from the ERA5 hourly 2-m air temperature series, resulting in a time series of $n=84$ block maxima. Such temperature series are frequently fitted using the GEV distribution, as for example reflected in work by the World Weather Attribution collaboration in estimating the anthropogenic contribution to the probability of observed heat extremes occurring [18, 19, 20].
The GEV and bGEV distributions were compared in terms of one-year-ahead probabilistic forecasts, using summed negative log likelihood (NLL) of the actual value in the forecast as the skill metric [21]. NLL, also referred to as the ignorance score [22], is closely related to cross-entropy and Kullback–Leibler divergence, information-theory measures of differences between a given distribution and an imperfect model of it [23, 24, 25], with lower NLL corresponding to closer agreement between the probabilistic forecast and the observed value. NLL has previously been used to quantitatively compare different methods for generating probabilistic temperature forecasts [6, 26, 27].
The probabilistic forecasts were generated by fitting, for each length $s$ from 30 up to $n-1$, maximum-likelihood GEV and bGEV distributions for the first $s$ annual maxima, and using those to generate a forecast for year $s+1$ that could be compared to the observed value. (The bGEV was allowed to have a shape parameter of either sign, using the extension outlined above.) NLL was therefore summed over 5400 separate forecasts (54 years for each of 100 locations).
@@
Figure 4. Example of a new sea level record from Hilo, Hawai’i in 2020, showing the histogram of the 1970-2019 annual maximum sea level and the forecasts for 2020 based on the GEV and bGEV. Both the GEV and bGEV appropriately shift the forecast probability distribution upward relative to the historic one due to global warming and have almost the same forecast mode, but the GEV distribution, with negative shape parameter, has zero probability mass past 0.9143 m, whereas the bGEV forecast has positive probability for the actual 0.916 m (vertical line).